Prepare sales and production budgets. A Sheville Company is preparing its master budget for 2022. Relevant...

  1. Prepare sales and production budgets.

A Sheville Company is preparing its master budget for 2022. Relevant data pertaining to its sales and production budgets are as follows.

  1. Sales. Sales for the year are expected to total 2,100,000 units. Quarterly sales, as a percentage of total sales, are 15%, 25%, 35%, and 25%, respectively. The sales price is expected to be $70 per unit for the first three quarters and $75 per unit beginning in the fourth quarter. Sales in the first quarter of 2023 are expected to be 10% higher than the budgeted sales volume for the first quarter of 2022.
  2. Production. Management desires to maintain ending finished goods inventories at 20% of the next quarter's budgeted sales volume.
  1. Prepare cash budget for a month.

The controller of Trenshaw Company wants to improve the company's control system by preparing a month-by-month cash budget. The following information is for the month ending July 31, 2022.

June 30, 2022, cash balance

$45,000

Dividends to be declared on July 15a

12,000

Cash expenditures to be paid in July for operating expenses

40,800

Amortization expense in July

4,500

Cash collections to be received in July

90,000

Merchandise purchases to be paid in cash in July

56,200

Equipment to be purchased for cash in July

20,000

Trenshaw Company wants to keep a minimum cash balance of $25,000.

  1. Prepare a cash budget for the month ended July 31, 2022, and indicate how much money, if any, Trenshaw Company will need to borrow to meet its minimum cash requirement.
  2. Explain how cash budgeting can reduce the cost of short-term borrowing.
Answer #1
Asheville Company
Sales Budget
For the Year Ending December 31, 2022
Quarter
1 2 3 4 Year
Expected unit sales          315,000.00                      525,000.00                      735,000.00         525,000.00         2,100,000.00
Unit selling price $               70.00 $                           70.00 $                           70.00 $              75.00
Total sales $ 22,050,000.00 $             36,750,000.00 $             51,450,000.00 $39,375,000.00 $ 149,625,000.00
Asheville Company
Production Budget
For the Year Ending December 31, 2022 Quarter
1 2 3 4 Year
Expected unit sales          315,000.00                      525,000.00                      735,000.00         525,000.00
Add: Desired ending finished goods units          105,000.00                      147,000.00                      105,000.00           69,300.00
Total required units          420,000.00                      672,000.00                      840,000.00         594,300.00
Less: Beginning finished goods units            63,000.00                      105,000.00                      147,000.00         105,000.00
Required production units          357,000.00                      567,000.00                      693,000.00         489,300.00         2,106,300.00

Quarter 2 =$F$12-25% 70 =C12C13 1 =$F$12.15% 70 =B12 B13 Year 2100000 3 =$F$12*35% 70 =D12 D13 4 =$F$12-25% 75 =E12E13 =SUM(B